Molarity(m)=Moles (mol) of solute/Liters (L) of solution

Molarity of standard NaOH from bottle = 0.1005 M

Molarity of NaOH added in titration (mL) = 10.47 mL

1. Moles of NaOH added in titration ______
Answer #1

Molarity(M)=Moles (mol) of solute/Liters (L)

It is the formula

We are given with

Molarity = 01005 M

Volume = 10.47 mL

Converting mL to L

10.47 mL\times \frac{1 L}{1000 mL}

=0.01047 L

Volume = 0.01047 L

Plug in the values in the formula

Molarity=\frac{mol}{volume in L}

0.1005 M=\frac{mol}{0.01045 L}

Multiply both sides with 0.01045 L

0.1005 \times 0.01045=\frac{mol}{0.01045}\times 0.01045

0.001050 = mol

Therefore, moles of NaOH in solution = 0.001050 mol = 1.050 x 10-3 mol